OpenStudy (anonymous):

Find the coordinates of the other endpoint when you are given the midpoint and one of the endpoints. P1 M P2 (12, 4) (2, 6) @timo86m help please!!:D

OpenStudy (anonymous):

mx=(x1+x2)/2 2=(12+x)/2 x=-8 6=(4+y)/2 y=8 so (-8,8) is p2